OpenStudy (anonymous):

The original value of a car is $18,000, and it depreciates (loses value) by 15% each year. What is the value of the car after three years?

OpenStudy (anonymous):

value after three years\[=18000\left( 1-\frac{ 15 }{100 } \right)^3=18000\left( \frac{ 85 }{100 } \right)^3=18000\left( \frac{ 17 }{20 } \right)^3=?\]